OpenStudy (anonymous):

Michael is 25 years younger than his father. The sum of their ages is 53. What is Michael’s age?

OpenStudy (lgbasallote):

let's say father's age = x since michael is 25 years younger than his father...that means michael's age is 25 less than his father so michael = x - 25 the sum of their age is 53 so.. (x) + (x - 25) = 53

OpenStudy (lgbasallote):

if you simplify that.. x + x - 25 = 53 2x - 25 = 53 addd 25 to both sides to isolate x 2x = 53 + 25 2x = 78 do you know the next step now?

OpenStudy (anonymous):

X = 39.. So, Michael's Age is : 39 - 25 = 14...

OpenStudy (lgbasallote):

alternate solution would be y = michael's age father's age will be y + 25 (because he's 25 years older) the sum is 53 so y + y + 25 = 53 2y + 25 = 53 2y = 53 - 25 2y = 28 this solution is to find michael's age...the one i did previously was father's age

OpenStudy (lgbasallote):

both just result the same thing